OpenStudy (anonymous):

how do i solve x/x^2-4 + 1/x+2=3

OpenStudy (anonymous):

x/x^2-4 + 1/x+2=3 x/(x-2)(x+2) + 1/x+2=3 x+x-2/(x-2)(x+2)=3 2x-2=3 *(x-2)(x+2) 2x-2=3(x^2 - 4) 3x^2 -2x-2=0 3x^2 -3x + x -2=0 3x(x -1) + 2(x -1)=0 (3x+2)((x - 1) =0 x= -2/3, 1 Hope I did not make any mistake ..let me know if thats correct or not?

OpenStudy (anonymous):

oh bad luck...found a mistake...thats the drawback of now writing but typing it

OpenStudy (anonymous):

let me correct it again

OpenStudy (anonymous):

3x^2 -2x-10=0 x= 2 +- sqrt (4+120) /6 x= 2+- 2 sqrt(31) /6 x= 1+- sqrt (31)/ 3 hope this time its correct...pls reply to let me know

OpenStudy (anonymous):

what is sqrt?

OpenStudy (anonymous):

square root

OpenStudy (anonymous):

well im completely confused because i dont know how to solve the problem..you lost me at the third step

OpenStudy (anonymous):

i took the lcm of (x-2)(x+2) and (x+2) which is (x-2)(x+2

OpenStudy (anonymous):

and shouldnt you have distributed the lcm throughout both sides?

OpenStudy (anonymous):

the lcm which we got , was multiplied with numerator on the right hand side which was 3

OpenStudy (anonymous):

fan me pls if you liked my help

OpenStudy (anonymous):

ok thanks..

OpenStudy (anonymous):

fan me plsssssssssss

OpenStudy (anonymous):

tnx ...

OpenStudy (anonymous):

welcum..

OpenStudy (anonymous):

gud luck with ur maths...